OpenStudy (amonoconnor):

Can someone help me check what I'm getting as the anti-derivative of "1/(3(x^(1/3)))"? Any and all help is greatly appreciated! :)

OpenStudy (anonymous):

that is really hard to read, but the idea is to write it in exponential form

OpenStudy (amonoconnor):

Yes, I apologize for the nasty way I had to type it! :/

OpenStudy (amonoconnor):

I got "(x^(4/3))/4" as the anti., but the answer key does NOT have that... Did I goof? I've checked my work like, literally 8 times!

OpenStudy (amonoconnor):

So... Here's my work, as I was taught to do theses :) 1/(3(x^(1/3))) = (1/3)(x^(1/3)) = (1/3)(x^((1/3) + 1))) = (1/3)(x^(4/3)) = [(1/3)/(4/3)]x^(4/3) = [(1/3)((3/4)]x^(4/3) = (3/12)x^(4/3) = (1/4)x^(4/3) = (x^(4/3))/4 + C

OpenStudy (anonymous):

It's x^(-1/3) not x^(1/3).

OpenStudy (anonymous):

The x is in the denominator.

OpenStudy (amonoconnor):

*rage* okay... Stupid me... I see how even though the 1/3 is factored, it's 1/(x^(1/3)) or x^(1/3), but... Yeah, just yeah. God I'm dumb. Thank you, oh amazing one. ;)

OpenStudy (amonoconnor):

Lols, I even forgot the negative there too. ;)